# The Reward Model Selection Crisis in Personalized Alignment

### v1

@@ 5 Pref-LaMP: A Benchmark for Direct Behavioral Evaluation
5.1 Dataset Construction
Pref-LaMP derives from LaMP-5 Salemi et al. (2024), pairing researchers’ abstracts with their titles. Both are author-written, capturing individual style. We construct preferences via hard negative mining: (1) encode abstracts with Qwen3-Embedding-0.6B, (2) retrieve top-$k$ similar abstracts from other users, (3) sample one retrieved abstract and use its title as $y^{(r)}$, (4) use original title as $y^{(c)}$. This ensures rejections are topically relevant but stylistically different. Human-written negatives avoid shortcut learning—initial LLM-generated rejections let linear probes detect generation artifacts rather than preference signals.
Pref-LaMP is the first benchmark enabling non-circular evaluation through user-authored completions, measurable via ROUGE and BERTScore.
5.2 Behavioral Alignment Metric
@@
Let $\mathcal{G}:\mathcal{X}\to\mathcal{Y}$ be a generation operator and $\mathcal{S}:\mathcal{Y}\times\mathcal{Y}\to[0,1]$ be a similarity measure. For user $k$ with test set $\mathcal{P}_{k}=\{(x_{j},y_{j}^{\text{GT}})\}_{j=1}^{M_{k}}$:
@@
We instantiate $\mathcal{G}$ with ARGS decoding (Eq. 1), and $\mathcal{S}$ with ROUGE-1 (lexical overlap), ROUGE-L (longest common subsequence), and BERTScore-F1 (semantic similarity). This measures whether generated outputs match what users actually write, providing ground-truth assessment of behavioral alignment.
@@ 6 Experimental Setup
6.1 Datasets and Models
TLDR Stiennon et al. (2020): Binary stylistic preferences, 10 training users (2,097 prefs/user), 31 adaptation users (100 prefs/user). PRISM Kirk et al. (2024): Pluralistic preferences, 1,232 training users (22.1 prefs/user), 139 adaptation users (14.5 prefs/user). Pref-LaMP (ours): User-authored completions, 485 training users (48.8 prefs/user), 126 adaptation users (49.2 prefs/user).
6.2 Models and Methods
@@
All reward models use LoRA rank 8, trained on $\mathcal{U}_{\text{train}}$ to learn shared $\theta$ and user-specific $\{z_{k}\}$. We evaluate six personalization methods: LoRE Bose et al. (2025) (learn reward bases and user-specific convex combination), LoRE-Alt (same as LoRE but alternates between bases and user specific parameter gradient steps), PReF Shenfeld et al. (2025) (collaborative filtering), PAL Chen et al. (2025a), VPL Poddar et al. (2024), MPU/MPU-Avg (a simple baselines of per-user MLPs), and P-DPO Li et al. (2024) (personalized direct preference optimization). Baselines include Global-RM (non-personalized Bradley-Terry using last token embedding as input to RM decoder), Global-RM-V2 (a sequence reward is average reward for all tokens), GenARM Xu et al. (2025) (autoregressive RM for token-level guidance), zero-shot generation, ICL (random demonstrations), and ICL-RAG (retrieved demonstrations).
6.3 Evaluation Protocol
We measure: (1) RM accuracy on adaptation/deployment users held-out preferences, (2) policy accuracy using prior/global/personalized scoring functions, (3) generation quality on Pref-LaMP via ROUGE-1/L and BERTScore against ground-truth, and (4) win rates where each method’s RM judges its own outputs versus zero-shot baseline.
7 Results Part 1: Discrimination Does Not Follow From Ranking
We first investigate whether reward model accuracy predicts discrimination ability under reward-guided decoding—i.e., whether methods that rank preferences accurately can also guide generation appropriately.
7.1 TLDR: Weak Correlation on Simple Data
@@
Table 2: Policy accuracy and reward model performance on TLDR + SmolLM. Base reward model is SmolLM2-135M
On TLDR’s binary preferences, most personalization methods (MPU, MPU-Avg, PReF) achieve policy accuracies below the prior baseline (62-65%), particularly at smaller scales. RM accuracy correlates weakly with policy accuracy (Kendall’s $\tau$ = 0.08–0.31), degrading as scale increases (Pearson $r$: 0.41 $\to$ 0.11 from 180M to 1.7B). PAL achieves highest RM accuracy (87.0%) but doesn’t support RGD, while LoRE—with lower RM accuracy (82.6%)—achieves superior policy accuracy ($\sim$81%). LoRE substantially outperforms all baselines with 16-20 point gains over prior policy, demonstrating genuine effectiveness and validating our evaluation framework.
7.2 PRISM: Personalization Fails on Pluralistic Data
@@
Table 3: Reward model accuracies of various algorithms using Qwen2.5-0.5B backbone on PRISM dataset. Global reward model V1 outperforms all other methods.
@@
Figure 1: RM vs Policy Accuracy Correlation on PRISM across model scales. Correlations remain consistently weak: Pearson $r$ ranges from 0.30–0.48, Spearman $\rho$ from 0.25–0.43, and Kendall $\tau$ from 0.17–0.29. While correlations slightly strengthen with scale, they remain far below what would be needed for RM accuracy to reliably predict policy performance. Notably, methods cluster vertically (similar RM accuracy, varying policy performance), demonstrating that high RM accuracy does not guarantee better policy understanding.
@@
Results are summarized in Table 3 and Figure 1. With reference to RM accuracies in Table 3, all personalized methods underperform Global RM (77.9% vs. 52.8-74.3%), with even the best (PReF, 74.3%) falling 3.6 points short. Meanwhile, RM-policy correlation, as shown in Fig 1, remains weak (Kendall’s $\tau$ = 0.17-0.29), slightly strengthening with scale ($r$: 0.31 $\to$ 0.48). Critically, VPL achieves highest policy accuracy (63.8-66.4%) despite 9.5 points lower RM accuracy than Global RM—a complete ranking inversion. Methods show minimal scaling gains, remaining in narrow bands (LoRE: 56.8-57.6%, VPL: 63.8-66.4%), suggesting limitations in learning from sparse pluralistic.
@@
Unlike TLDR/SmolLM2 where correlations degraded with scale ($r$ = 0.41→0.11), PRISM/Qwen2.5 shows strengthening correlations ($r$ = 0.31→0.48). Whether this reflects dataset differences, model architecture, or their interaction remains unclear. Regardless, even at 7B, correlations remain too weak for reliable prediction.
Reconciling with Prior Claims. LoRE’s released code fails to reproduce claimed gains over Global-RM [^1] https://github.com/facebookresearch/LoRe/issues/1; our reimplementation of their linear probe setup yields almost identical accuracy (us: 55.59% vs. theirs: 55.55%) validating our imlpementation. PReF was evaluated on PERSONA, not PRISM which they admittedly fail on (Appendix D). PAL and VPL used different datasets (TLDR, non-overlapping tasks).
Implication: RM accuracy cannot guide deployment—methods with 10+ point gaps perform identically or invert in ranking. This implies that RM accuracy does not measure behaviorally-relevant properties for policy guidance. Reward models can rank preferences accurately while failing to provide useful token-level guidance.
@@
Recommendation. Evaluations must assess downstream policy understanding across multiple datasets and scales—not just RM accuracy. The RM-policy disconnect demands new metrics measuring reward models’ suitability for guiding generation, rather than pairwise ranking accuracy alone.
8 Results Part 2: Discrimination Does Not Imply Generation
@@
Given the weak correlation between RM accuracy and policy discrimination ability under RGD, we now ask: even when methods achieve high policy accuracy—demonstrably preferring chosen over rejected responses—do they generate outputs that behaviorally align with user preferences?
8.1 Pref-LaMP: Complete Decoupling Revealed
LoRE LoRE-Alt MPU MPU-Avg PAL PReF VPL Vanilla Vanilla-v2 GenARM
RM Acc $65.60$ $\pm 7.89$ $84.96$ $\pm 0.48$ $65.26$ $\pm 1.03$ $67.30$ $\pm 0.30$ $53.77$ $\pm 0.28$ $51.46$ $\pm 3.84$ $43.63$ $\pm 3.77$ $84.96$ $\pm 0.13$ $84.69$ $\pm 0.07$ $83.89$ $\pm 0.60$
Table 4: Reward model accuracies using Qwen2.5-0.5B on Pref-LaMP. Vanilla RMs again outperform personalization methods.
@@ 8.1 Pref-LaMP: Complete Decoupling Revealed
Global RMs dominate (84.7-85.0%), with GenARM (83.9%) and LoRE-Alt close. Most personalization methods fail (VPL: 43.6%, PReF: 51.5%). Figure 2 reveals three critical failures:
(1) Architecture provides no advantage. GenARM, designed for token-level guidance, performs comparably to vanilla methods despite similar RM accuracy.
(2) RM accuracy decouples from generation. Best reward methods (LoRE, Vanilla RM, MPU-Avg) achieve identical ROUGE-1 (45.9-46.0) and BERTScore (70.5-70.8) at 7B despite 20-point RM accuracy gaps. High RM accuracy guarantees nothing about generation quality.
(3) Reward guidance fails at scale. At 0.5-1.5B, reward methods improve over zero-shot. At 3-7B, ICL-RAG dominates, achieving 48.9 ROUGE-1 and 72.2 BERTScore at 7B—3.0 and 1.5 point gaps over best reward methods. RGD becomes counterproductive as base capability increases.
@@
GenARM 100.0% 100.0% 100.0% 99.8%
LoRE 73.7% 72.9% 73.8% 73.7%
@@
Vanilla RM 99.9% 99.9% 99.9% 99.9%
Vanilla RM-v2 98.4% 97.2% 89.4% 84.4%
Table 5: Win rate: fraction of examples where each RM judges its own RGD output as better than zero-shot. High win rates reveal severe reward hacking—methods claim near-perfect improvement despite ground-truth metrics showing minimal or negative gains.
Win rates expose reward hacking: GenARM claims 100% improvement over zero-shot despite ROUGE-1 being marginally worse. Vanilla RM claims 99.9% superiority while producing outputs identical to methods 20 points lower in RM accuracy. Circular evaluation actively misleads—methods optimize their scoring functions then declare success while actual alignment remains unmeasured.
Figure 3: ROUGE-1 vs. number of in-context examples. ICL-RAG consistently outperforms random ICL, with performance scaling with both model size and context length.
ICL-RAG improves steadily with demonstrations and scale, reaching $\sim$49 ROUGE-1 at 7B with 5 shots. Larger models show no saturation, effectively leveraging context without reward guidance. This reveals RGD’s failure is not inherent to personalization but to its operationalization mechanism.
8.2 No Metric Predicts Generation Quality
@@
Table 6 exposes the complete failure of standard evaluation: RM accuracy, policy accuracy, and win rates show negligible to negative correlations with generation quality (Kendall’s $\tau$ = -0.188 to -0.017). Even policy accuracy—measuring discrimination under the actual deployment mechanism—cannot predict behavioral alignment. Win rates are often most negative, confirming reward hacking.
@@
The takeaway is stark: no existing metric predicts whether methods will generate aligned outputs. Ground-truth evaluation on user-authored completions is not optional—it’s necessary.
9 Discussion and Recommendations
Our findings reveal a fundamental evaluation crisis: RM accuracy (Kendall’s $\tau$ = 0.08-0.31 with policy performance) and discrimination ability both fail to predict generation quality. Methods with 20-point RM accuracy gaps produce identical outputs, while reward models claiming 99% win rates show no ground-truth improvement.
For practitioners: Use ICL-RAG for models $\geq$3B. Reward modeling adds complexity without benefit at scale.
For researchers: (1) Report generation quality on ground-truth completions, not just RM accuracy. (2) Evaluate across model scales to detect scale-dependent failures. (3) Compare against ICL baselines. (4) Build behavioral benchmarks like Pref-LaMP with user-authored completions.
The discrimination-generation gap reflects RGD’s nature as a complex dynamical system. At each generation step, two forces interact: base model logits and reward scores. These can agree or conflict, with dominance shifting across token positions. Policy accuracy measures whether the scoring function on average prefers tokens in chosen responses, but generation is sensitive to outliers—a single poorly-scored token can derail trajectories (causing length changes, topic drift). Moreover, discrimination and generation are distinct objectives: successfully ranking complete sequences doesn’t guarantee steering generation away from undesirable regions of the output space. Understanding these dynamics is critical future work.
Limitations: We focus on inference-time methods (RGD) because it represents the only scalable deployment path—per-user RL fine-tuning remains computationally infeasible for realistic user populations. While RL-based personalization might exhibit different failure modes, it cannot address the fundamental problem we expose: practitioners need to select reward models for deployment, and current metrics fail to predict performance under the only viable deployment mechanism. Generalization beyond our three datasets requires investigation across diverse personalization dimensions (dialogue, values, humor). Our theoretical explanations of the discrimination-generation gap remain preliminary and warrant deeper mechanistic study.
10 Conclusion
We have demonstrated that personalized alignment from preference data faces a fundamental evaluation crisis: standard reward model accuracy metrics do not predict deployment performance under reward-guided decoding. Through systematic evaluation across three datasets and four model scales, we show that (1) RM accuracy correlates only weakly with discrimination ability, (2) discrimination ability doesn’t predict generation quality, and (3) simple in-context learning dominates reward-guided methods for models $\geq$3B parameters.
These findings expose that the field has been optimizing proxy metrics divorced from actual deployment objectives. Current personalized alignment methods fundamentally fail to operationalize user preferences into behavioral adaptation under realistic deployment constraints. The introduction of Pref-LaMP—the first benchmark with ground-truth user completions—enables non-circular evaluation and reveals the extent of this failure through direct behavioral assessment.
For practitioners, the path forward is clear: use ICL with retrieval for personalized generation at scale. For researchers, our work highlights the urgent need for ground-truth behavioral evaluation, theoretical understanding of the discrimination-generation gap, and development of methods that can effectively translate preferences into generation guidance. Only by measuring what actually matters—whether models generate outputs aligned with user preferences—can the field make progress on truly personalized alignment.

### v2

@@ 5 Pref-LaMP: A Benchmark for Direct Behavioral Evaluation
Dataset ConstructionPref-LaMP derives from LaMP-5 Salemi et al. (2024), pairing researchers’ abstracts with their titles. Both are author-written, capturing individual style. We construct preferences via hard negative mining: (1) encode abstracts with Qwen3-Embedding-0.6B, (2) retrieve top-$k$ similar abstracts, (3) sample one retrieved abstract as $x$ and use its title as $y^{(l)}$, (4) use original title as $y^{(w)}$. This ensures rejections are topically relevant but different in title formulation [^1] Human-written rather than LLM-written negatives avoid shortcut learning. Initial LLM-generated rejections let linear probes detect generation artifacts rather than preference signals..
Pref-LaMP is the first benchmark enabling direct behavioural evaluation of personalization through user-authored completions, measurable via ROUGE and BERTScore.
Behavioral Alignment Metric  We evaluate end-to-end behavioural alignment by comparing user-generated and personalized model responses.
@@
Let $\mathcal{G}:\mathcal{X}\to\mathcal{Y}$ be a generation operator and $\mathcal{S}:\mathcal{Y}\times\mathcal{Y}\to\mathcal{R}$ be a similarity measure. For user $k$ with test set $\mathcal{P}_{k}=\{(x_{j},y_{j}^{\text{GT}})\}_{j=1}^{M_{k}}$:
@@
We instantiate $\mathcal{G}$ with ARGS decoding (Eq. 1), zero-shot generation and ICL generation. Meanwhile, $\mathcal{S}$ in instantiated with ROUGE-1 (lexical overlap), ROUGE-L (longest common subsequence), and BERTScore-F1 (semantic similarity). This measures whether generated outputs match what users actually write, providing ground-truth assessment of behavioral alignment.
@@ 6 Experimental Setup
Datasets and Models  We consider datasets: TLDR Stiennon et al. (2020): Binary stylistic preferences, 10 training users (2,097 prefs/user), 31 adaptation users (100 prefs/user). PRISM Kirk et al. (2024): Pluralistic preferences, 1,232 training users (22.1 prefs/user), 139 adaptation users (14.5 prefs/user). Pref-LaMP (ours): User-authored completions, 485 training users (48.8 prefs/user), 126 adaptation users (49.2 prefs/user).
@@
Models and Methods  All reward models use LoRA rank 8, trained on $\mathcal{U}_{\text{train}}$ to learn shared $\theta$ and user-specific $\{z_{k}\}$. We evaluate six personalization methods: LoRE Bose et al. (2025) (learn reward bases and user-specific convex combination), LoRE-Alt (same as LoRE but alternates between bases and user specific parameter gradient steps), PReF Shenfeld et al. (2025) (collaborative filtering), PAL Chen et al. (2025a), VPL Poddar et al. (2024), MPU/MPU-Avg (a simple baselines of per-user MLPs), and P-DPO Li et al. (2024) (personalized direct preference optimization). Baselines include Global-RM (non-personalized Bradley-Terry using last token embedding as input to RM decoder), Global-RM-V2 (a sequence reward is average reward for all tokens), GenARM Xu et al. (2025) (autoregressive RM for token-level guidance), zero-shot generation, ICL (random demonstrations), and ICL-RAG (retrieved demonstrations).
Evaluation Protocol  We measure: (1) RM accuracy on adaptation users’ held-out preferences (Eq. 2), (2) Adaption users’ policy accuracy vs prior (no-reward) and global reward baselines (Eq. 3), (3) generation quality on Pref-LaMP via ROUGE-1/L and BERTScore against ground-truth (Eq. 8), and (4) win rates where each method’s RM judges its own outputs versus zero-shot baseline.
7 Result 1: Reward Model Accuracy Does Not Imply Policy Accuracy
We first investigate whether reward model accuracy predicts policy accuracy under reward-guided decoding. I.e., whether personal rewards that rank preferences well can guide policies to do the same.
@@ 7 Result 1: Reward Model Accuracy Does Not Imply Policy Accuracy
Table 2: Policy accuracy and reward model performance on TLDR + SmolLM. Base reward model is SmolLM2-135M. Top: Personal rewards. Mid: Global alignment.
TLDR: Weak Correlation on Simple Data  We first evaluate the popular TDLR dataset’s simple binary style preferences in Table 2. The main observation is that upstream RM accuracy correlates weakly with downstream policy accuracy (Kendall’s $\tau$ = 0.08–0.31), degrading as scale increases (Pearson $r$: 0.41 $\to$ 0.11 from 180M to 1.7B).
Additionally, most personalization methods (MPU, MPU-Avg, PReF) fail to adapt policies, achieving policy accuracies below the prior baseline (62-65%), particularly at smaller scales. PAL achieves highest RM accuracy (87.0%) but doesn’t support RGD, while LoRE—with lower RM accuracy (82.6%)—achieves superior policy accuracy ($\sim$81%). Overall, LoRE substantially outperforms all competitors as well as prior and global reward policies, demonstrating genuine effectiveness for inference-time adaptation and validating our evaluation framework.
Personalized Global
@@
Table 3: Reward model accuracies of various personal alignment methods using Qwen2.5-0.5B backbone on PRISM dataset. All personal alignment methods underperform non-personal Global Reward Model V1.
@@
Figure 1: RM vs Policy Accuracy Correlation on PRISM across model scales. Correlations remain consistently weak: Pearson $r$ ranges from 0.30–0.48, Spearman $\rho$ from 0.25–0.43, and Kendall $\tau$ from 0.17–0.29. While correlations slightly strengthen with scale, they remain far below what would be needed for RM accuracy to reliably predict policy performance. Notably, methods with similar RM accuracy can have substantially different policy accuracy and vice-versa, demonstrating that the standard RM metric is not a reliable proxy for deployment performance.
@@
PRISM: Personalization Fails on Pluralistic Data  Results for the more complex PRISM benchmark are summarized in Table 3 and Figure 1. All personalization methods fail in terms of their policy accuracy underperforming the global RM baseline (77.9% vs. 52.8-74.3%) in Table 3. Meanwhile, RM-policy correlation, as shown in Fig 1, remains weak (Kendall’s $\tau$ = 0.17-0.29), slightly strengthening with scale ($r$: 0.31 $\to$ 0.48). Critically, VPL achieves highest policy accuracy (63.8-66.4%) despite 9.5 points lower RM accuracy than Global RM—a complete ranking inversion.
@@
In terms of scale, methods show minimal scaling gains, remaining in narrow bands (LoRE-Alt: 57.1-58.5%, VPL: 63.8-66.4%). Unlike TLDR/SmolLM2 where correlations degraded with scale ($r$ = 0.41→0.11), PRISM/Qwen2.5 shows strengthening correlations ($r$ = 0.31→0.48). Whether this reflects dataset differences, model architecture, or their interaction remains unclear. Regardless, even at 7B, correlations remain too weak for choosing methods based on RM accuracy.
Discussion  Our careful control evaluation shows wide failure of prior personal alignment methods both in terms of beating global alignment baselines, and in terms of the standard metric of RM accuracy not corresponding to downstream policy accuracy. We attribute this to a mixture of released code not reproducing results, missing non-personal baselines, and inconsistent non-comparable choice of datasets in prior evaluations. See Appendix D and E for further discussion.
Implication: Personal RM accuracy does not reflect performance during policy inference and cannot guide choice of reward model for deployment: Methods with 10+ point RM gaps can perform identically as adapted policies; methods with near identical RM accuracy can have 10+ point gaps in policy accuracy; and alignment methods can invert in ranking between reward and policy evaluations.
@@
Recommendation. Future personal alignment methods must specify a policy adaptation strategy, and assess downstream policy understanding across multiple datasets and scales—not just upstream personal reward accuracy. The RM-policy disconnect demands new metrics measuring reward models’ suitability for guiding generation, rather than pairwise ranking accuracy alone.
8 Result 2: Preference Discrimination Does Not Imply Generation Quality
@@
Given the weak correlation between RM accuracy and policy discrimination ability under RGD, we now ask: even when methods achieve high policy accuracy—demonstrably preferring chosen over rejected responses—do they actually generate outputs that behaviorally align with user preferences?
8.1 Pref-LaMP: Complete Decoupling
Global Personalized
Acc GenARM Global RM Global RM-v2 LoRE LoRE-Alt MPU MPU-Avg PAL PReF VPL ICL
RM $83.89^{\pm 0.60}$ $84.96^{\pm 0.13}$ $84.69^{\pm 0.07}$ $65.60^{\pm 7.89}$ $84.96^{\pm 0.48}$ $65.26^{\pm 1.03}$ $67.30^{\pm 0.30}$ $53.77^{\pm 0.28}$ $51.46^{\pm 3.84}$ $43.63^{\pm 3.77}$ -
Policy $71.34^{\pm 0.56}$ $68.28^{\pm 0.00}$ $78.02^{\pm 0.94}$ $69.65^{\pm 0.31}$ $65.76^{\pm 6.60}$ $63.46^{\pm 8.11}$ $63.16^{\pm 6.70}$ - $8.79^{\pm 0.70} 66$$6.75^{\pm 2.33}$ $75.67^{\pm 0.44}$
Table 4: Pref-LaMP preference ranking accuracy for RMs and adapted policies (Qwen2.5-3B). RM personalization does not clearly outperform global RMs for either upstream RM or downstream policy preference ranking.
@@ 8.1 Pref-LaMP: Complete Decoupling
We first study Pref-LaMP with preference ranking evaluation in Table 4. They key observation is that for this challenging task, similarly to PRISM (Table 3), personal alignment methods struggle to surpass Global RM baselines – for both the standard proxy metric of upstream RM accuracy, as well as our downstream policy accuracy. Only LoRE-Alt come close to the global baselines in RM accuracy.
We next move to analysing behavioural generation quality of the policies – as uniquely enabled by our Pref-LAMP dataset, in Figure 2 and raw results in Appendix C. We see that: (1) The top personal alignment methods for upstream ranking accuracy (LoRE-Alt in Table 4) tend to underperform in downstream generation quality. (2) A few personal alignment methods can surpass the zero-shot baseline but produce comparative performance to Global RM baseline. However, the better methods for generation (e.g, MPU-Avg and LoRe) are worse for upstream ranking (Table 4). Both these observations reflect decoupling between upstream RM accuracy and downstream behavioural generation. This shows that downstream generation quality evaluation is a crucial missing component of standard evaluation practice.
Our end-to-end behavioural evaluation also allows direct comparison between existing RM-focused personal alignment approaches and ICL. From Table 4, we can see that ICL actually achieves better policy preference ranking than the personal RMs. In terms of generation quality, Figure 2 shows that direct application of ICL surpasses both the baselines and prior personal alignment methods at 3-7B scale. This suggests that practitioners today should use simple ICL in favour of complex RM-based alignment approaches.
@@
Global GenARM 100.0% 100.0% 100.0% 99.8%
Global RM 99.9% 99.9% 99.9% 99.9%
Global RM-v2 98.4% 97.2% 89.4% 84.4%
Personalized LoRE 73.7% 72.9% 73.8% 73.7%
@@
Table 5: Win rate: fraction of examples where each RM judges its own RGD output as better than zero-shot. High win rates reveal severe reward hacking—methods claim near-perfect improvement despite ground-truth metrics showing minimal or negative gains (Fig 2).
Discussion: Reconciling Prior Claims  How can we reconcile prior papers’ claims of successful RM+RGD based non-personal alignment with the often negative results from our experiments? The evaluation protocol of prior RGD-based analyses involved guiding generation with a RM, and then evaluating the resulting generations using the same RM Khanov et al. (2024). The issue with win-rates scored in this way is circularity. If RGD-adaptation can hack the RM (find a ’false positive’ response that the RM accepts, while not actually reflecting user preferences), using the same RM to evaluate the result produces overly optimistic results.
To study this protocol, we report win Rate vs zero-shot in Table 5, which confirms the risk of ‘circular’ evaluation. Using the same RM to guide decoding and judge completions vs a baseline suffers from reward hacking/overoptimization El-Mhamdi and Hoang (2024). GenARM claims 100% improvement over zero-shot despite ROUGE-1 being marginally worse. Global RM claims 99.9% superiority while producing outputs identical to methods 20 points lower in RM accuracy. The circular evaluation protocol is thus vulnerable to reward hacking, and only appears to solve personal alignment.
Despite their limitations, ranking held-out preferences (Table 2,4) does not suffer from this – because the RM is not used to generate; neither does our ground-truth evaluation (Figure. 2) – because the RM generation is compared to ground-truth.
Additional Analysis: ICL  We provide further analysis in Appendix C.1 showing ICL and ICL-RAG improve further with shots higher than 8 (as used in the main text).
@@
No Metric Predicts Generation Quality  To summarise, we considered the standard RM accuracy, RGD-win rate, and our policy-accuracy metrics, all of which are discriminative ranking metrics. Table 6 correlates each of these against our end-to-end generation quality metric. All exhibit negligible to negative correlations with generation quality (Kendall’s $\tau$ = -0.188 to -0.017), with the negative correlations suggesting reward hacking.
@@
Takeaway: No existing metric predicts whether personalization methods will generate aligned outputs. Ground-truth evaluation on user-authored completions is a necessary evaluation component.
9 Discussion and Conclusion
Our findings reveal an evaluation crisis in personal alignment research: RM accuracy is uncorrelated with policy accuracy ($\tau$ = 0.08-0.31), and method rankings can completely invert between upstream and downstream evaluations. Using Pref-LaMP—the first benchmark with ground-truth user completions, we show discriminative metrics fail to predict generation quality (Table 6): reward models claiming 99% win rates show no improvement over baselines in ground-truth similarity. The field has been optimizing proxy metrics divorced from deployment objectives.
On a more positive note, we highlight that in contrast to these issues with personal rewards and their evaluations, simple in-context learning dominates reward-guided methods for models $\geq$3B parameters, while being easy and reliable to implement.
Practical Recommendations.  Practitioners should use ICL with retrieval for 3B+ models; reward modeling adds complexity without benefit at scale. Researchers should: (1) evaluate complete pipelines end-to-end, not just reward model accuracy, (2) include policy accuracy and ground-truth behavioral metrics, (3) test across model scales to detect scale-dependent effects, (4) build behavioral benchmarks with user-authored completions and (5) compare against ICL baselines and focus future research effort on developing such amortized approaches to personal alignment.
9.1 Limitations
We focus on RGD because it represents a key scalable deployment path—per-user RL fine-tuning remains computationally infeasible for realistic populations. A fundamental challenge with RGD is that it assumes reward models can be token-wise factorized to provide local guidance at each generation step, which is a known source of error when this assumption is violated Li et al. (2025). While GenARM is specifically designed to address this limitation through token-level autoregressive reward training, it still exhibits the same performance gaps we observe across other methods. This suggests the problem runs deeper than factorization alone—the disconnect between preference learning and generation guidance may be fundamental to the inference-time adaptation paradigm.
Our use of three datasets goes beyond most prior work, which often used only one or contrived datasets. However, our results do show some facets of dataset dependence, so future work should aim to establish larger multi-dataset benchmark suits to thoroughly test personalization across more dimensions of interest.